f(x)=secx²+3x²+yx²
1
Expert's answer
2020-02-07T12:16:38-0500

"f(x,y)=2xsec(x^2)+3x^2+yx^2"

"\frac{\partial f(x,y)}{\partial x}=2x\frac{sin(x^2)}{cos^2(x^2)}+x(6+2y)"

"\frac{\partial f(x,y)}{\partial y}=x^2"